A 0.25 ?? object moves due east at 2.1 ?/?. What force is needed to cause...

  1. A 0.25 ?? object moves due east at 2.1 ?/?. What force is needed to cause it to move due north at 3.6 ?/? in 1.52 ??

  2. An object of mass 2.8 ?? moves at 1.1 ?/? in the +? direction. If an impulse of 1.3 ? ∙ ? ?3 is applied, what is its final momentum?
Answer #1

change in momentum (Ap) = P-P; apmus-mii - m (JfJ7) = 0-25kg (3.6 m/s 1 - 2.1m/s 7) Ap=(-0.5251 +0.99) kgomis TAPI = 0.525) +